# NVM-in-Cache: Repurposing Commodity 6T SRAM Cache into NVM Analog Processing-in-Memory Engine using a Novel Compute-on-Powerline Scheme

### v1

@@ Abstract
The rapid growth of deep neural network (DNN) workloads has significantly increased the demand for large-capacity on-chip SRAM in machine learning (ML) applications, with SRAM arrays now occupying a substantial fraction of the total die area. To address the dual challenges of storage density and computation efficiency, this paper proposes an NVM-in-Cache architecture that integrates resistive RAM (RRAM) devices into a conventional 6T-SRAM cell, forming a compact 6T-2R bit-cell. This hybrid cell enables Processing-in-Memory (PIM) mode, which performs massively parallel multiply-and-accumulate (MAC) operations directly on cache power lines while preserving stored cache data. By exploiting the intrinsic properties of the 6T-2R structure, the architecture achieves additional storage capability, high computational throughput without any bit-cell area overhead. Circuit- and array-level simulations in GlobalFoundries 22nm FDSOI technology demonstrate that the proposed design achieves a throughput of 0.4 TOPS and 491.78 TOPS/W. For 128 row-parallel operations, the CIFAR-10 classification is demonstrated by mapping a Resnet-18 neural network, achieving an accuracy of 91.27%. These results highlight the potential of the NVM-in-Cache approach to serve as a scalable, energy-efficient computing method by re-purposing existing 6T SRAM cache architecture for next-generation AI accelerators and general purpose processors.
Index Terms:
Analog in-memory computing, SRAM, 6T-Cell, NVM-in-cache, processing in-memory (PIM).
@@ I Introduction
The growing demand for data-intensive computations, coupled with transistor scaling in accordance with the Moore’s law and the increasing complexity of consumer workloads, has driven steady growth in on-chip cache capacity over the years [1]. With the emergence of AI and deep neural networks, this trend has become even more critical to achieving state-of-the-art performance and precision. As the size of these networks increases, recent processors have required larger on-chip caches to meet performance needs [2]. Although several studies have explored cache architectures using 7T [3], 8T [4], 9T [5], and 10T [6] SRAM cells, the conventional 6T SRAM still remains almost exclusively used in commercial processors due to its high density and fast access speed [7]. However, with the rapid growth in computational workloads and the slowing benefits of technology scaling, there is a need to explore alternative memory technologies. Non-volatile memories (NVMs) offer smaller cell footprints, enabling higher density, and exhibit significantly lower leakage compared to conventional SRAM, which suffers from substantial leakage power due to short channel effects at advanced technology nodes [8]. Some NVM options are Resistive RAM (RRAM) [9], Magnetic RAM (MRAM) [10], and Phase Change Memory (PCM) [11]. Despite their advantages, NVMs generally suffer from higher write latency and limited endurance, increased write power consumption, and lower signal-to-noise ratio (SNR) compared to SRAM [12]. Thereby, despite their density and power benefits, they are not suitable for on-chip high-speed cache applications. This motivates the integration of 6T SRAM with NVM devices, effectively combining the advantages of both NVM and 6T SRAM inside on-chip cache hierarchies, without incurring substantial (almost no) bit-cell area overhead.
@@
In this work, we present an NVM-in-Cache architecture which integrates NVM within the 6T SRAM cells, thereby repurposing 6T SRAM to perform massively parallel analog processing-in-memory (PIM) operations using a novel compute-on-powerline scheme. Continuous technology scaling has enabled an increase in the number of processor cores on a chip, but memory bandwidth and energy efficiency have not scaled proportionally, leading to the well-known memory wall [13]. In deep learning applications, the growing data transfer between memory and the processor has become a major bottleneck for energy-efficient computation. This has motivated the concept of PIM, also referred to as Computing-in-Memory (CIM), where computation is performed within the memory array to minimize data movement and improve efficiency in AI workloads. Prior NVM-based PIM designs have demonstrated this concept but often suffer from limited SNR ratio [14], low ON/OFF conductance ratio [15], and fabrication challenges for large-scale integration [16]. Among different NVM technologies, RRAM is notable for its relatively high conductance ratio [17], and has also been demonstrated as a scalable, high-density memory in commercial fabrication processes [18]. In this work we use RRAM for design and analysis, but the presented scheme is applicable to other NVMs as well. Note, while NVMs address some of the limitations of 6T SRAM, it is not an ideal direct replacement for cache memory due to its high write latency and energy consumption. This motivates the integration of NVM (RRAM) with conventional cache architectures, enabling efficient in-memory computation without compromising cache functionality.
@@
Previous work on SRAM-based PIM [19, 20, 21] has shown promising results but is generally not well-suited for cache architectures, as many designs employ more than six transistors per bit-cell, significantly reducing cache density. Some studies [22, 23] have demonstrated PIM using standard 6T SRAM cells; however, these approaches require the cache to flush its existing data in order to store the neural network weights for in-memory computation. After computation, the cache data must be reloaded, which introduces additional latency and increases energy consumption due to extra data movement.
@@ II-A Resistive Random Access Memory (RRAM) Device
RRAM is a type of resistive non-volatile memory that stores information based on the resistance state of the device with a high conductance ratio. It typically consists of silver (Ag) and platinum (Pt) electrodes separated by a conductive filament. The bipolar RRAM device operates with a positive SET voltage to switch to a low-resistance state (LRS) and a negative RESET voltage to return to a high-resistance state (HRS). The device switches from high- to low-resistance upon exceeding the SET voltage and returns to high-resistance when the RESET voltage is reached. A compact Verilog-A model that captures the non-linearity of the device and key parameters has been employed [24]. The LRS and HRS states represent the binary values 1 and 0, which are used to store neural network weights in our proposed scheme. These weights can be represented in an analog form either by physically adjusting the dimensions of the RRAM device or by tuning the programming states to achieve different resistance levels [18]. The latter approach is more flexible than altering the physical structure. The capability of RRAM is discussed in detail in Section 4.
@@ II-B Cache Architecture
We provide a brief overview of the last-level cache (LLC) architecture by referencing the Intel multi-core Xeon and AMD Zen processor [25, 26, 27, 28] as shown in Fig. 1. The shared LLC is accessed by the multi-core processor via a ring/mesh interconnect. [25] architecture consists of 8 LLC slices, each comprising a 20-way, 2.5 MB cache divided into 80 banks of 32 KB each. In addition to the data banks, each slice contains the corresponding 1-way tag, cache-valid (CV) bits, state, and least recently used (LRU) structures. Each 32 KB bank is further divided into 16 KB subarrays, and each subarray consists of two 8 KB (256 $\times$ 256) SRAM arrays, along with the necessary decoders, peripheral circuitry, and input/output interfaces. The focus of this work is to retain the overall architecture while replacing the conventional 6T SRAM cells with a 6T-2R structure. This modification not only preserves the standard cache functionality but also enables PIM capabilities, allowing computations to be performed directly within the memory arrays. By doing so, the proposed design leverages the benefits of NVM without compromising the cache’s original performance and organization.
@@ III Design of the Proposed 6T-2R Bit-cell
In this section, we describe the proposed 6T-2R bitcell in detail. A conventional 6T SRAM layout consists of ground lines (VSS), wordlines (WL), and specifically isolated power lines (VDD) which provides an opportunity to integrate non-volatile memory (NVM) between pull up devices and isolated VDD power lines. In our design the wordlines are separated (WL1 and WL2) which does not incur any additional area overhead [29]. Further, two RRAM devices, $R_{LEFT}$ and $R_{RIGHT}$, are integrated along the VDD lines, while the VSS lines are extended across each row of the cache array to form a gated-GND structure [30] as shown in Fig. 2. This arrangement allows the 6T SRAM cell to be augmented with non-volatile memory devices in a seamless manner, thereby doubling the effective storage capacity without incurring any bit-cell area overhead. Beyond storage enhancement, the proposed structure also enables efficient support for PIM operations. The gated-GND control provides additional flexibility in managing the current flow through the cell, which is critical for ensuring reliable computation while retaining stored volatile data in the 6T cell. The bitcell itself is organized symmetrically, where the left and right halves are defined by powerlines (VDD1/VDD2), wordlines (WL1/WL2), bitlines (BL/BLB), and a dedicated shared control signal (V1/V2) that governs the gated-GND operation.
@@ III-A NVM Programming
During the programming stage, only the 2T-2R portion of the 6T-2R bitcell, as highlighted in Fig. 3, is utilized. Lets assume, the initial resistance state of the RRAM devices is set to HRS. To switch to LRS, a SET voltage must be applied, whereas a RESET voltage is required to transition back from HRS to LRS. The programming phase leverages the concept of wordline overdrive[31]. Programming from HRS to LRS requires two cycles, one for $R_{LEFT}$ and one for $R_{RIGHT}$, with each cycle having a pulse width of 4 ns. In the first cycle, $R_{LEFT}$ is programmed by overdriving WL1 and WL2 to 2 V, applying 2 V to BL and 0 V to BLB, and setting the gated-GND control signal (V1 and V2) to 0 V. Since the required SET voltage across the RRAM is 1.2 V, both VDD1 and VDD2 are grounded during this operation. In the second cycle, $R_{RIGHT}$ is programmed using the same method, except with BL = 0 V and BLB = 2 V, while the remaining signals are unchanged. During the LRS programming phase, the bitlines must be driven in complementary fashion to selectively activate the PMOS access transistors (M2 for $R_{LEFT}$ and M4 for $R_{RIGHT}$). This ensures a proper current path is established, enabling reliable programming of the RRAM devices.
@@
Programming the RRAM devices to the HRS requires only a single cycle. During this operation, both wordlines (WL1 and WL2) are overdriven to 2 V, while BL and BLB are held at 0 V. The supply lines VDD1 and VDD2 are biased at 2 V, and the gated-GND control signal (V1 and V2) is set to 0 V. Unlike LRS programming, HRS programming does not require separate cycles for $R_{LEFT}$ and $R_{RIGHT}$, since both can be programmed in parallel. This is because, with BL and BLB grounded, both storage nodes (Q and QB) are also forced to 0 V, thereby turning on the PMOS transistors (M2 and M4) and enabling current flow from VDD1/VDD2 to BL/BLB.
@@
To validate successful programming, a read operation is performed immediately after the programming phase in a single cycle. For the read, VDD1 and VDD2 are held at VDD (0.8), WL1 and WL2 are biased at VDD, and BL and BLB are used to measure the current for 1ns. The current is then measured at the bitlines: if the RRAM device is in LRS, the measured current is high, whereas for HRS the current drops to low value, thereby confirming the programmed state. In our work, we use a programming voltage of 2 V, which is relatively high but determined by the specific RRAM model employed in this work. Use of RRAM devices optimized for low programming voltage will result in lesser overdrive. It is also important to note that programming is destructive to the SRAM data; however, in most neural network–based PIM applications, particularly for inference, the frequency of reads far outweighs programming operations, minimizing latency and power overhead. Furthermore, both $R_{LEFT}$ and $R_{RIGHT}$ are programmed to the same state (LRS or HRS) during PIM, which preserves the inherent symmetry of the SRAM cell.
@@ III-B SRAM Hold/Read/Write Mode
The proposed 6T-2R structure preserves the conventional read, write, and hold functionality of a standard 6T SRAM cache bit-cell. During the hold and read operation, the intrinsic advantage of placing the RRAM devices along the power lines becomes evident. Importantly, the resistance state of the RRAMs has no impact on data retention as shown in Fig. 4. In the hold state, the supply voltages VDD1 and VDD2 are maintained at 0.8 V, which corresponds to the nominal operating voltage of the cell, while both wordlines (WL1 and WL2) are held low (0 V), V1 and V2 is set to 0.8 V as in a conventional 6T SRAM. Consider the case where the storage nodes Q/QB hold the values 1/0. Under this condition, PMOS transistor M2 remains ON, while M4 remains OFF. On the left half of the cell, the potential at node Q is equal to 0.8 V. Since there is no voltage difference between VDD1 (0.8 V) and node Q (0.8 V), no current flows through the path, thereby ensuring that Q remains strongly held at logic ‘1’ irrespective of the RRAM resistance. By the cross-coupled inverter structure, QB is correspondingly maintained at logic ‘0’. A similar argument applies when Q/QB = 0/1, in which case the complementary transistors switch roles to preserve the state. This demonstrates that the addition of RRAM devices along the power lines does not interfere with the intrinsic bistability of the SRAM latch, and the stored data is reliably retained regardless of the resistance state of the RRAMs. The read and write operations of the proposed 6T-2R bitcell, when operating in SRAM mode, are identical to those of a conventional 6T SRAM. All control signals, including the wordlines (WL1 and WL2), bitlines (BL/BLB), and supply voltages, function in the same manner as in a standard SRAM cell [31].
Figure 5: PIM dot-product operation in the proposed 6T-2R bitcell: (a) The operation is performed on the left half, where the weight is stored in $R_{LEFT}$ and the IA is applied to WL1, while the right half (QB) retains the cache data; (b) The same operation is performed on the right half; (c) Table illustrating the cases for computing the dot product in two phases.

### v2

@@ Abstract
The rapid growth of deep neural network (DNN) workloads has significantly increased the demand for large-capacity on-chip SRAM in machine learning (ML) applications, with SRAM arrays now occupying a substantial fraction of the total die area. To address the dual challenges of storage density and computation efficiency, this paper proposes an NVM-in-Cache architecture that integrates resistive RAM (RRAM) devices into a conventional 6T-SRAM cell, forming a compact 6T-2R bit-cell. This hybrid cell enables Processing-in-Memory (PIM) mode, which performs massively parallel multiply-and-accumulate (MAC) operations directly on cache power lines while preserving stored cache data. By exploiting the intrinsic properties of the 6T-2R structure, the architecture achieves additional storage capability, high computational throughput without any bit-cell area overhead. Circuit- and array-level simulations in GlobalFoundries 22nm FDSOI technology demonstrate that the proposed design achieves a throughput of 0.4 TOPS and 452.34 TOPS/W. For 128 row-parallel operations, the CIFAR-10 classification is demonstrated by mapping a Resnet-18 neural network, achieving an accuracy of 91.76%. These results highlight the potential of the NVM-in-Cache approach to serve as a scalable, energy-efficient computing method by re-purposing existing 6T SRAM cache architecture for next-generation AI accelerators and general purpose processors.
@@ I Introduction
The growing demand for data-intensive computations, coupled with transistor scaling in accordance with the Moore’s law and the increasing complexity of consumer workloads, has driven steady growth in on-chip cache capacity over the years [cachesizetrend]. With the emergence of AI and deep neural networks, this trend has become even more critical to achieving state-of-the-art performance and precision. As the size of these networks increases, recent processors have required larger on-chip caches to meet performance needs [burd2022zen3]. Although several studies have explored cache architectures using 7T [7tsram], 8T [8tsram], 9T [9tsram], and 10T [10tsram] SRAM cells, the conventional 6T SRAM still remains almost exclusively used in commercial processors due to its high density and fast access speed [tsmcsram]. However, with the rapid growth in computational workloads and the slowing benefits of technology scaling, there is a need to explore alternative memory technologies. Non-volatile memories (NVMs) offer smaller cell footprints, enabling higher density, and exhibit significantly lower leakage compared to conventional SRAM, which suffers from substantial leakage power due to short channel effects at advanced technology nodes [sramleak]. Some NVM options are Resistive RAM (RRAM) [rram], Magnetic RAM (MRAM) [mram], and Phase Change Memory (PCM) [pcm]. Despite their advantages, NVMs generally suffer from higher write latency and limited endurance, increased write power consumption, and lower signal-to-noise ratio (SNR) compared to SRAM [nvmprob]. Thereby, despite their density and power benefits, they are not suitable for on-chip high-speed cache applications. This motivates the integration of 6T SRAM with NVM devices, effectively combining the advantages of both NVM and 6T SRAM inside on-chip cache hierarchies, without incurring substantial (almost no) bit-cell area overhead.
@@
In this work, we present an NVM-in-Cache architecture which integrates NVM within the 6T SRAM cells, thereby repurposing 6T SRAM to perform massively parallel analog processing-in-memory (PIM) operations using a novel compute-on-powerline scheme. Continuous technology scaling has enabled an increase in the number of processor cores on a chip, but memory bandwidth and energy efficiency have not scaled proportionally, leading to the well-known memory wall [memorywall]. In deep learning applications, the growing data transfer between memory and the processor has become a major bottleneck for energy-efficient computation. This has motivated the concept of PIM, also referred to as Computing-in-Memory (CIM), where computation is performed within the memory array to minimize data movement and improve efficiency in AI workloads. Prior NVM-based PIM designs have demonstrated this concept but often suffer from limited SNR ratio [snrprob], low ON/OFF conductance ratio [lowonoff], and fabrication challenges for large-scale integration [fabricationchallenge]. Among different NVM technologies, RRAM is notable for its relatively high conductance ratio [rram_ratio], and has also been demonstrated as a scalable, high-density memory in commercial fabrication processes [rramden]. In this work we use RRAM for design and analysis, but the presented scheme is applicable to other NVMs as well. Note, while NVMs address some of the limitations of 6T SRAM, it is not an ideal direct replacement for cache memory due to its high write latency and energy consumption. This motivates the integration of NVM (RRAM) with conventional cache architectures, enabling efficient in-memory computation without compromising cache functionality.
@@
Previous work on SRAM-based PIM [srampim1, srampim2, srampim3] has shown promising results but is generally not well-suited for cache architectures, as many designs employ more than six transistors per bit-cell, significantly reducing cache density. Some studies [6tsrampim1, 6tsrampim2] have demonstrated PIM using standard 6T SRAM cells; however, these approaches require the cache to flush its existing data in order to store the neural network weights for in-memory computation. After computation, the cache data must be reloaded, which introduces additional latency and increases energy consumption due to extra data movement.
@@ II-A Resistive Random Access Memory (RRAM) Device
RRAM is a type of resistive non-volatile memory that stores information based on the resistance state of the device with a high conductance ratio. It typically consists of silver (Ag) and platinum (Pt) electrodes separated by a conductive filament. The bipolar RRAM device operates with a positive SET voltage to switch to a low-resistance state (LRS) and a negative RESET voltage to return to a high-resistance state (HRS). The device switches from high- to low-resistance upon exceeding the SET voltage and returns to high-resistance when the RESET voltage is reached. A compact Verilog-A model that captures the non-linearity of the device and key parameters has been employed [rrammodel]. The LRS and HRS states represent the binary values 1 and 0, which are used to store neural network weights in our proposed scheme. These weights can be represented in an analog form either by physically adjusting the dimensions of the RRAM device or by tuning the programming states to achieve different resistance levels [rramden]. The latter approach is more flexible than altering the physical structure. The capability of RRAM is discussed in detail in Section 4.
@@ II-B Cache Architecture
We provide a brief overview of the last-level cache (LLC) architecture by referencing the Intel multi-core Xeon and AMD Zen processor [intel1, intel2, neuralcache, amdzen5] as shown in Fig. 1. The shared LLC is accessed by the multi-core processor via a ring/mesh interconnect. [intel1] architecture consists of 8 LLC slices, each comprising a 20-way, 2.5 MB cache divided into 80 banks of 32 KB each. In addition to the data banks, each slice contains the corresponding 1-way tag, cache-valid (CV) bits, state, and least recently used (LRU) structures. Each 32 KB bank is further divided into 16 KB subarrays, and each subarray consists of two 8 KB (256 $\times$ 256) SRAM arrays, along with the necessary decoders, peripheral circuitry, and input/output interfaces. The focus of this work is to retain the overall architecture while replacing the conventional 6T SRAM cells with a 6T-2R structure. This modification not only preserves the standard cache functionality but also enables PIM capabilities, allowing computations to be performed directly within the memory arrays. By doing so, the proposed design leverages the benefits of NVM without compromising the cache’s original performance and organization.
@@ III Design of the Proposed 6T-2R Bit-cell
In this section, we describe the proposed 6T-2R bitcell in detail. A conventional 6T SRAM layout consists of ground lines (VSS), wordlines (WL), and specifically isolated power lines (VDD) which provides an opportunity to integrate non-volatile memory (NVM) between pull up devices and isolated VDD power lines. In our design the wordlines are separated (WL1 and WL2) which does not incur any additional area overhead [wlsep]. Further, two RRAM devices, $R_{LEFT}$ and $R_{RIGHT}$, are integrated along the VDD lines, while the VSS lines are extended across each row of the cache array to form a gated-GND structure [gatedvdd] as shown in Fig. 2. This arrangement allows the 6T SRAM cell to be augmented with non-volatile memory devices in a seamless manner, thereby doubling the effective storage capacity without incurring any bit-cell area overhead. Beyond storage enhancement, the proposed structure also enables efficient support for PIM operations. The gated-GND control provides additional flexibility in managing the current flow through the cell, which is critical for ensuring reliable computation while retaining stored volatile data in the 6T cell. The bitcell itself is organized symmetrically, where the left and right halves are defined by powerlines (VDD1/VDD2), wordlines (WL1/WL2), bitlines (BL/BLB), and a dedicated shared control signal (V1/V2) that governs the gated-GND operation.
@@ III-A NVM Programming
During the programming stage, only the 2T-2R portion of the 6T-2R bitcell, as highlighted in Fig. 3, is utilized. Lets assume, the initial resistance state of the RRAM devices is set to HRS. To switch to LRS, a SET voltage must be applied, whereas a RESET voltage is required to transition back from HRS to LRS. The programming phase leverages the concept of wordline overdrive[assisttech]. Programming from HRS to LRS requires two cycles, one for $R_{LEFT}$ and one for $R_{RIGHT}$, with each cycle having a pulse width of 4 ns. In the first cycle, $R_{LEFT}$ is programmed by overdriving WL1 and WL2 to 2 V, applying 2 V to BL and 0 V to BLB, and setting the gated-GND control signal (V1 and V2) to 0 V. Since the required SET voltage across the RRAM is 1.2 V, both VDD1 and VDD2 are grounded during this operation. In the second cycle, $R_{RIGHT}$ is programmed using the same method, except with BL = 0 V and BLB = 2 V, while the remaining signals are unchanged. During the LRS programming phase, the bitlines must be driven in complementary fashion to selectively activate the PMOS access transistors (M2 for $R_{LEFT}$ and M4 for $R_{RIGHT}$). This ensures a proper current path is established, enabling reliable programming of the RRAM devices.
@@
Programming the RRAM devices to the HRS requires only a single cycle. During this operation, both wordlines (WL1 and WL2) are overdriven to 2 V, while BL and BLB are held at 0 V. The supply lines VDD1 and VDD2 are biased at 2 V, and the gated-GND control signal (V1 and V2) is set to 0 V. Unlike LRS programming, HRS programming does not require separate cycles for $R_{LEFT}$ and $R_{RIGHT}$, since both can be programmed in parallel. This is because, with BL and BLB grounded, both storage nodes (Q and QB) are also forced closed to 0 V, thereby turning on the PMOS transistors (M2 and M4) and enabling current flow from VDD1/VDD2 to BL/BLB.
@@
To validate successful programming, a read operation is performed immediately after the programming phase in a single cycle. For the read, BL and BLB are held at VDD (0.8 V), WL1 and WL2 are biased at VDD depending on whether to read $R_{LEFT}$ and $R_{RIGHT}$, and VDD1 and VDD2 are used to measure the current for 1ns. The current is then measured at the powerlines: if the RRAM device is in LRS, the measured current is high, whereas for HRS the current drops to low value, thereby confirming the programmed state. In our work, we use a programming voltage of 2 V, which is relatively high but determined by the specific RRAM model employed in this work. Use of RRAM devices optimized for low programming voltage will result in lesser overdrive. It is also important to note that programming is destructive to the SRAM data; however, in most neural network–based PIM applications, particularly for inference, the frequency of reads far outweighs programming operations, minimizing latency and power overhead. Furthermore, both $R_{LEFT}$ and $R_{RIGHT}$ are programmed to the same state (LRS or HRS) during PIM, which preserves the inherent symmetry of the SRAM cell.
@@ III-B SRAM Hold/Read/Write Mode
The proposed 6T-2R structure preserves the conventional read, write, and hold functionality of a standard 6T SRAM cache bit-cell. During the hold and read operation, the intrinsic advantage of placing the RRAM devices along the power lines becomes evident. Importantly, the resistance state of the RRAMs has no impact on data retention as shown in Fig. 4. In the hold state, the supply voltages VDD1 and VDD2 are maintained at 0.8 V, which corresponds to the nominal operating voltage of the cell, while both wordlines (WL1 and WL2) are held low (0 V), V1 and V2 is set to 0.8 V as in a conventional 6T SRAM. Consider the case where the storage nodes Q/QB hold the values 1/0. Under this condition, PMOS transistor M2 remains ON, while M4 remains OFF. On the left half of the cell, the potential at node Q is equal to 0.8 V. Since there is no voltage difference between VDD1 (0.8 V) and node Q (0.8 V), no current flows through the path, thereby ensuring that Q remains strongly held at logic ‘1’ irrespective of the RRAM resistance. By the cross-coupled inverter structure, QB is correspondingly maintained at logic ‘0’. A similar argument applies when Q/QB = 0/1, in which case the complementary transistors switch roles to preserve the state. The hold operation is also tested on a 128 row 6T-2R array. This demonstrates that the addition of RRAM devices along the power lines does not interfere with the intrinsic bistability of the SRAM latch, and the stored data is reliably retained regardless of the resistance state of the RRAMs. The read and write operations of the proposed 6T-2R bitcell, when operating in SRAM mode, are identical to those of a conventional 6T SRAM. All control signals, including the wordlines (WL1 and WL2), bitlines (BL/BLB), and supply voltages, function in the same manner as in a standard SRAM cell [assisttech].
Figure 5: PIM dot-product operation in the proposed 6T-2R bitcell: (a) The operation is performed on the left half, where the weight is stored in $R_{LEFT}$ and the IA is applied to WL1, while the right half (QB) retains the cache data; (b) The same operation is performed on the right half. (c)-(d) Transient behaviour of the circuit for cache data Q = 1 and QB = 0.